# Correlated helimagnetic configuration in a nonsymmorphic magnetic nodal semimetal

### v2

@@ Abstract
Nonsymmorphic magnetic Weyl semimetal materials such as ReAlX (Re=rare earth, X=Si/Ge) provide a unique opportunity to explore the correlated phenomena between Weyl fermions and nontrivial magnetic configurations. To be specific, we study a lattice model in which the magnetic configuration is determined by the competition among ferromagnetic (FM) interaction, the Dzyaloshinsky-Moriya interaction, and the Kondo coupling $K_{0}$ to the Weyl fermion. Both quantum and finite-temperature phase transitions between FM and correlated nesting helical configurations are found. Different from the uncorrelated helimagnet that decouples from the Weyl fermions, this correlated helimagnet induces a magnetic Brillouin zone with a $K_{0}$-dependent nesting in the band structure of the conduction electrons instead of the monopole-like Weyl cone. By measuring the current induced by the chiral magnetic effect on the conduction electron with nesting Weyl nodes, one can distinguish the correlated nesting helical order from the ferromagnetism because the chiral magnetic effect is considerably suppressed in the former case. These properties we find here may explain the experimental observations in ReAlX.
Introduction. Recently, the nodal semimetals in a set of nonsymmorphic magnetic materials, ReAlX(Re=rare earth, X=Si/Ge), attract widespread attention Xu2017 ; Chang2018 ; Sanchez2020 ; Destraz2020 and a type of Weyl semimetals with breaking both the time-reversal $\Theta$ and inversion $P$ symmetries was found in ReAlX Tafti2023 ; Tafti2021 ; Tafti2020 ; Broholm2021 ; Li2023 ; Alam2023 ; Yu2023 ; Yang2023 ; Terashima2024 ; Yamada2024 . These magnetic materials belong to the body-centered tetragonal crystal system with a non-centrosymmetric space group $I4_{1}md$. Due to the $P$ breaking, the Weyl nodes possess different energies. Meanwhile, the $f$-orbital electrons of the rare earth elements provide a strong magnetic moment, so that a strong Ruderman-Kittel-Kasuya-Yosida (RKKY) interaction can exist by mediating Weyl fermions. The RKKY interaction induces competition among the Heisenberg, Kitaev-Ising, and Dzyaloshinsky-Moriya (DM) interactions Xiao2015 ; ZhouPRB2015 ; ZhouPRB2017 , which leads to the emergence of several nontrivial magnetic configurations, such as spiral, helical, hedgehog, skyrmion, and meron-antimeron pairs Tretiakov2021 .
@@
Physically, those nontrivial magnetic orders rotate the Weyl fermion’s spin configuration in real space via the Kondo coupling, which competes with the monopole topology of the Weyl nodes in momentum space. This competition may further induce new correlated phenomena and correlated topological states of matter that lack clear explanations yet.
@@
In this Letter, we focus on the correlated nesting helical magnetic configuration (CNHMC), which has been observed in ReAlX materials. For example, since the ratio between out-of-plane and in-plane magnetic susceptibility $\chi_{c}/\chi_{a}$ is very close in SmAlSi at $2K$, a helimagnet with ${\bf Q}=\frac{2\pi}{3}(0.979,0.979,0)$ emerges, which is consistent with the nesting vector between Weyl nodes Tafti2023 . In CeAlGe, a phase transition from the canted magnetic order to helimagnet was observed at $4.5K$. The helimagnet also provides a nesting vector between the Weyl nodes and stabilizes magnetic fluctuations from the magnetic transition temperature to the Curie temperature Yu2023 . This phase transition to the helimagnet and the CNHMC effect are also reported for PrAlSi Tafti2020 , NdAlSi Broholm2021 ; Yamada2024 , CeAlSi Tafti2021 , and NdAlGe Yang2023 .
@@
However, the mechanism of the magnetic phase transition and the origin of the emergence of the CNHMC in those materials remain unclear. The specific mechanism behind the anomalous transport phenomena observed near $T=12.8$K in CeAlGe Yu2023 , such as the nearly zero magnetoresistance, the switch from negative to positive magnetoresistance, and the maximal suppression of thermal conductivity by magnetic fields is not yet fully understood. In Ref. Yu2023 , in which two of the present authors joined, a phenomenological analysis of these properties was partially done by assuming uncorrelated Weyl node nesting, but studies on the CNHMC and the magnetic phase transition are lacking.
@@
Conventionally, the magnetic configuration ${\bf M}$ and Weyl fermions are separately studied. However, in ReAlX, the strong Kondo coupling plays an important role, so this separated treatment is not valid. In this Letter, we find that the strong coupling problem is exactly solvable at the Fermi surface. We then take this correlated ${\bf M}$ configuration as a background field to deal with the conduction electrons. When we determine the phase diagram of the correlated magnetism, we fix the fermion spin configuration. With these approximations, we see the phase transitions from an ordered phase, e.g., the ferromagnetic (FM) one, to the CNHMC both at zero temperature and finite temperature. We calculate the current induced by the chiral magnetic effect (CME) both in the FM phase and the CNHMC and find that the current for the latter is an order of magnitude less than that for the former. This may explain the nearly zero magnetoresistance near $T=12.8$K in CeAlGe Yu2023 .
@@
Lattice model. The common model study for the materials ReAlX starts from the calculation of the band structure of this body-centered tetragonal crystal system with a non-centrosymmetric space group $I4_{1}md(109)$. The Weyl nodes are found near the Fermi level and are coupled to the local moment from the $f$-electron in the rare-earth element. Such a strategy is fine but difficult to understand the basic physics due to the complicated numerics.
@@
Due to the crystal symmetry and observed large anomalous Hall effects Tafti2020 ; Tafti2021 ; Yang2023 ; Li2023 ; Terashima2024 , the conduction electrons in ReAlX are Weyl semimetals with both $\Theta$ and $P$ symmetry breaking. To see the basic physics behind the numerics, we study a simplified cubic lattice model in which the conduction electrons are Weyl semimetals without $\Theta$ and $P$ symmetries and have two Weyl nodes that are adjusted to be near the Fermi level. The electrons couple to ${\bf M}_{i}$ via the Kondo coupling while the nearest neighbor ${\bf M}_{i}$ and ${\bf M}_{j}$ are coupled through the FM Heisenberg and the DM interactions:
@@
For simplicity, we restrict $H_{W}$ to be a minimal two-band model feiye2016 ; Trivedi2017 and the Weyl nodes are located at $(0,0,\pm k_{0})$ with $\cos k_{0}=(m/2-t_{x}-t_{y})/t_{z}$. $c_{j}^{\dagger}$ (the spin index omitted) is the creation operator of the conduction electron at site $j$ whose spin ${\bf s}_{j}=\frac{1}{2}c_{j}^{\dagger}{\boldsymbol{\sigma}}c_{j}$. For the FM case, $J_{0}<0$. In the following, we use $t_{x}=t_{y}=t_{z}=t^{\prime}_{x}=t^{\prime}_{y}=t=1,m=5t$ and $k_{0}=\pi/3$. The chemical potential has been set to zero. $H_{K}$ is the on-site Kondo coupling between ${\bf s}_{i}$ and ${\bf M}_{i}$ with the coupling strength $K_{0}$. $H_{H}$ and $H_{DM}$ are the nearest-neighbor FM Heisenberg and DM interactions, respectively.
When solving the conduction electron problem, we use the adiabatic approximation, i.e., assuming the conduction electrons are fast-moving with respect to the fluctuations of ${\bf M}$ when ${\bf M}$ is considered a background. On the other hand, the configuration of $\bf M$ is determined by the competition of the variational ground state energies when the conduction electron spin is assumed to be polarized by ${\bf M}$ ’s configuration, the CNHMC or FM.
@@
where $\uparrow$ and $\downarrow$ are the spin indices. The last term in Eq. (Correlated helimagnetic configuration in a nonsymmorphic magnetic nodal semimetal) is the Kondo coupling between the Weyl fermion and the helical magnetic configuration. Notice that the momentum shift from $k_{z}$ to $k_{z}\pm Q$ comes from the helical wave number $Q$.
Continuum limit and CNHMC effects. Before going to the numerical calculations for the lattice model, we consider the continuum limit of Eq. (Correlated helimagnetic configuration in a nonsymmorphic magnetic nodal semimetal) to show that the correlated helimagnet is an allowed magnetic ground state configuration. When $|Q|\ll 2|k_{0}|$, due to the energy offset $b_{0}$ between the two Weyl nodes in a non-centrosymmetric system, we can integrate out one Weyl fermion that is below the chemical potential, say, $\psi_{-}$. The effective equations of motion of ${\bf M}$ and the conduction electron $\psi=\psi_{+}$ at a given Weyl valley in the continuum limit are given by Yu16 ; suppl1
@@
where $J\propto K_{0}^{2}-2J_{0}$ and $D\propto v_{-}K_{0}^{2}+4D_{0}$ rkky1 ; rkky2 . In the correlated case, the helimagnet remains an exact $E_{e}=0$ solution with $Q=\frac{D\pm\sqrt{D^{2}+16JK_{0}}}{4J},$ which reduces to the uncorrelated helimagnet or the FM when $K_{0}=0$ i.e., either $Q=|{\bf Q}|=|D/2J|$ and or $Q=0$ Nagaosa2012 . When $K_{0}\neq 0$, ${\bf M}$ and $\psi$ are correlated via the Kondo coupling. The energy spectrum of the conduction electron $\psi$ is shown in Fig. 1. The zero mode solutions are localized at the boundaries of the magnetic Brillouin zone which is introduced due to the helimagnet (For details, see suppl1 ).
Figure 1: (color online) The energy spectrum in the continuum limit for Weyl fermions coupled with a helimagnet along the $q_{z}$-direction. We set $Q=K_{0}=1$. The helimagnet introduces a magnetic Brillouin zone. The red points stand for the Weyl nodes and the blue points stand for the zero modes (see suppl1 ).
@@
and $\Phi=(c_{+\uparrow},c_{+\downarrow},c_{-\uparrow},c_{-\downarrow})$ with $\pm$ labeling two Weyl valleys; $\sigma$ and $\tau$ are the matrices that act on spin and valley indices, respectively. $k_{x}^{\prime}=2t_{x}^{\prime}k_{x}$, $k_{y}^{\prime}=2t_{y}^{\prime}k_{y}$, $k_{z}^{\prime}=2t_{z}(\sin{k_{0}})k_{z}$, $a_{0}={v_{0}(\cot{k_{0}})}/{t}$, and $b_{0}=2v_{0}(\sin{k_{0}})$. The spectra of Hamiltonian (4) read $E_{\pm,\pm}=a_{0}k^{\prime}_{z}\\ \pm\sqrt{{\bf k}^{\prime 2}+(K_{0}/2)^{2}+b_{0}^{2}\pm\sqrt{K_{0}^{2}(k_{y}^{\prime 2}+k_{z}^{\prime 2})+4b_{0}^{2}{\bf k}^{\prime 2}}}$. In the $K_{0}\rightarrow 0$ limit, $b_{0}$ becomes the energy offset between the two Weyl nodes, and $a_{0}$ controls the type of the Weyl node.
Band structures of the fermion. We now turn back to the lattice model. Similar to dealing with integer quantum Hall effects with the lattice model Kohmoto1989 ; Kohmoto1990 ; Hatsugai1993 (also see suppl2 for a brief review), one defines $|\Psi\rangle=\sum_{j=1}^{q}\psi_{j}c^{\dagger}(k_{x},k_{y},k_{z}^{0}+Qj)|0\rangle,$ where $Q=2\pi p/q$ with $p$ and $q$ coprime, $k_{z}=k_{z}^{0}+Qj$, and $\psi_{j+q}=\psi_{j}$. Then, the Schrödinger equation $H_{0}|\Psi\rangle=E|\Psi\rangle$ is reduced to the Harper equations
[[[[ $m-2t_{x}\cos k_{x}-2t_{y}\cos k_{y}-2t_{z}\cos(k_{z}^{0}+Qj)]\sigma_{z}\psi_{j}$ (5)

### v3

@@ Abstract
Nonsymmorphic magnetic Weyl semimetal materials such as ReAlX (Re=rare earth, X=Si/Ge) provide a unique opportunity to explore the correlated phenomena between Weyl fermions and nontrivial magnetic configurations. To be specific, we study a lattice model in which the magnetic configuration is determined by the competition among ferromagnetic (FM) interaction, the Dzyaloshinskii-Moriya interaction, and the Kondo coupling $K_{0}$ to the Weyl fermion. Both quantum and finite-temperature phase transitions between FM and correlated nesting helical configurations are found. Different from the uncorrelated helimagnet that decouples from the Weyl fermions, this correlated helimagnet induces a magnetic Brillouin zone with a $K_{0}$-dependent nesting in the band structure of the conduction electrons instead of the monopole-like Weyl cone. By measuring the current induced by the chiral magnetic effect on the conduction electron with nesting Weyl nodes, one can distinguish the correlated nesting helical order from the ferromagnetism because the chiral magnetic effect is considerably suppressed in the former case. These properties we find here may explain the experimental observations in ReAlX.
I Introduction
Recently, the nodal semimetals in a set of nonsymmorphic magnetic materials, ReAlX(Re=rare earth, X=Si/Ge), attract widespread attention Xu2017 ; Chang2018 ; Sanchez2020 ; Destraz2020 and a type of Weyl semimetals with breaking of both the time-reversal $\Theta$ and inversion $P$ symmetries was found in ReAlX Tafti2023 ; Tafti2021 ; Tafti2020 ; Broholm2021 ; Li2023 ; Alam2023 ; Yu2023 ; Yang2023 ; Terashima2024 ; Yamada2024 . These magnetic materials belong to the body-centered tetragonal crystal system with a non-centrosymmetric space group $I4_{1}md$. Due to the $P$ breaking, the Weyl nodes possess different energies. Meanwhile, the $f$-orbital electrons of the rare earth elements provide a strong magnetic moment, so that a strong Ruderman-Kittel-Kasuya-Yosida (RKKY) interaction can exist by mediating Weyl fermions. The RKKY interaction induces competition among the Heisenberg, Kitaev-Ising, and Dzyaloshinskii-Moriya (DM) interactions Xiao2015 ; rkky2 ; ZhouPRB2015 ; ZhouPRB2017 ; LL2025 , which leads to the emergence of several nontrivial magnetic configurations, such as spiral, helical, hedgehog, skyrmion, and meron-antimeron pairs Tretiakov2021 .
@@ I Introduction
Physically, those nontrivial magnetic orders rotate the Weyl fermion’s spin configuration in real space via the Kondo coupling, which competes with the monopole topology of the Weyl nodes in momentum space. This competition may further induce new correlated phenomena and correlated topological states of matter that lack clear explanations thus far.
@@
In this paper, we focus on the correlated nesting helical magnetic configuration (CNHMC), which has been observed in ReAlX materials. For example, since the ratio between out-of-plane and in-plane magnetic susceptibility $\chi_{c}/\chi_{a}$ is very close in SmAlSi at $2K$, a helimagnet with ${\bf Q}=\frac{2\pi}{3}(0.979,0.979,0)$ emerges, which is consistent with the nesting vector between Weyl nodes Tafti2023 . In CeAlGe, a phase transition from the canted magnetic order to the helimagnet was observed at $4.5K$. The helimagnet also provides a nesting vector between the Weyl nodes and stabilizes magnetic fluctuations from the magnetic transition temperature to the Curie temperature Yu2023 . This phase transition to the helimagnet and the CNHMC effect are also reported for PrAlSi Tafti2020 , NdAlSi Broholm2021 ; Yamada2024 , CeAlSi Tafti2021 , and NdAlGe Yang2023 .
@@
However, the mechanism of the magnetic phase transition and the origin of the emergence of the CNHMC in those materials remain unclear. The specific mechanism behind the anomalous transport phenomena observed near T=12.8T=12.8K in CeAlGe Yu2023 , such as the nearly zero magnetoresistance, the switch from negative to positive magnetoresistance, and the maximal suppression of thermal conductivity by magnetic fields is not yet fully understood. In Ref. Yu2023 , in which two of the present authors participated, a phenomenological analysis was conducted by treating the magnetic order and Weyl fermions separately (termed uncorrelated Weyl node nesting), while the CNHMC effects and magnetic phase transition remain unexplored.
@@
Conventionally, the magnetic configuration $\bf M$} and Weyl fermions are separately studied. However, in ReAlX, the strong Kondo coupling plays an important role, so this separated treatment is not valid. In this paper, we develop an exact solution to the Kondo coupling problem in the continuum limit, identifying a characteristic zero-energy mode at the Fermi surface associated with helical magnetic ordering. We then take this correlated $\bf M$} configuration as a background field to deal with the conduction electrons. When we determine the phase diagram of the correlated magnetism, we fix the fermion spin configuration. With these approximations, we see the phase transitions from an ordered phase, e.g., the ferromagnetic (FM) one, to the CNHMC both at zero temperature and finite temperature. We calculate the current induced by the chiral magnetic effect (CME) both in the FM phase and the CNHMC and find that the current for the latter is an order of magnitude less than that for the former. This may explain the nearly zero magnetoresistance near T=12.8T=12.8K in CeAlGe Yu2023 .
The rest of this paper is organized as follows. In Sec. II, we construct a minimal two-band model for a Weyl semimetal that breaks both time-reversal and inversion symmetries and is coupled to the magnetic configuration $\bf M$}. We also consider the continuum limit of the model with the helical magnetic configuration, and an exact solution is found at zero energy. In Sec. III, we draw the band structures of the lattice model with CNHMC and find that the band structure forms a magnetic Brillouin zone, and the Weyl nodes are folded onto each other for small Kondo coupling. For large Kondo coupling, the Weyl nodes move towards each other, eventually annihilating and opening a gap. In this section, we also consider the magnetic phase diagram by calculating the ground state energy per site at zero temperature and the free energy per site at finite temperature. We find a competition between the ferromagnetic phase and the helical one, which is consistent with the experiments. In Sec. IV, we calculate the CME of the lattice model. We find that the corresponding magnetoresistance of CNHMC is strongly suppressed compared to that of the ferromagnetic case, which may relate to the nearly zero magnetoresistance observed in CeAlGe Yu2023 . The final part is devoted to conclusions.
II Lattice model and continuum limit
II.1 Lattice model
@@
The common model study for the materials ReAlX starts from the calculation of the band structure of this body-centered tetragonal crystal system with a non-centrosymmetric space group $I4_{1}md(109)$. The Weyl nodes are found near the Fermi level and are coupled to the local moment from the $f$ electron in the rare-earth element. Such a strategy is fine but it is difficult to understand the basic physics due to the complicated numerics.
@@
Due to the crystal symmetry and observed large anomalous Hall effects Tafti2020 ; Tafti2021 ; Yang2023 ; Li2023 ; Terashima2024 , the conduction electrons in ReAlX are Weyl semimetals with both $\Theta$ and $P$ symmetry breaking. Remarkably, ReAlX-type materials exhibit dramatic changes in transport properties (e.g., magnetoresistance) across the phase transition to the helimagnetically ordered state. To see the basic physics behind the numerics, we study a simplified cubic lattice model in which the conduction electrons are Weyl semimetals without $\Theta$ and $P$ symmetries and have two Weyl nodes that are adjusted to be near the Fermi level. The electrons couple to ${\bf M}_{i}$ via the Kondo coupling while the nearest neighbor ${\bf M}_{i}$ and ${\bf M}_{j}$ are coupled through the FM Heisenberg and the DM interactions:
@@
[…]
Beyond the helical magnetic order, other complex magnetic configurations (e.g., canted orders) and physical phenomena emerge due to multiple Weyl node pairs and lattice non-symmorphic features, though these are beyond our present scope. Our work concentrates on the ferromagnetic-to-helical phase transition and its topological characteristics, while a four-band model addressing multiple Weyl pairs will be developed in future studies.
When solving the conduction electron problem, we use the adiabatic approximation, i.e., assuming the conduction electrons are fast-moving with respect to the fluctuations of $\bf M$} when $\bf M$} is considered a background. On the other hand, the configuration of $\bf M$ is determined by the competition of the variational ground state energies when the conduction electron spin is assumed to be polarized by $\bf M$} ’s configuration, the CNHMC or ferromagnet.
@@
where $\uparrow$ and $\downarrow$ are the spin indices. The last term in Eq. (II.1) is the Kondo coupling between the Weyl fermion and the helical magnetic configuration. Notice that the momentum shift from $k_{z}$ to $k_{z}\pm Q$ comes from the helical wave number $Q$.
II.2 Continuum limit and CNHMC effects
To demonstrate the energetic preference for helical ordering in Kondo-coupled Weyl systems, we first establish its stability in the continuum limit [Eq. (II.1)] before validating with lattice simulations, and demonstrate that the correlated helimagnet constitutes an allowed ground-state configuration. When $|Q|\ll 2|k_{0}|$, due to the energy offset $b_{0}$ between the two Weyl nodes in a non-centrosymmetric system, we can integrate out one Weyl fermion that is below the chemical potential, say, $\psi_{-}$. The effective equations of motion of $\bf M$} and the conduction electron $\psi=\psi_{+}$ at a given Weyl valley in the continuum limit are given by Yu16
@@ II.2 Continuum limit and CNHMC effects
where $J\propto K_{0}^{2}-2J_{0}$ and $D\propto v_{-}K_{0}^{2}+4D_{0}$ Xiao2015 ; rkky2 . In the correlated case, the helimagnet remains an exact $E_{e}=0$ solution with $Q=\frac{D\pm\sqrt{D^{2}+16|J|K_{0}}}{4|J|},$ which reduces to the uncorrelated helimagnet or the ferromagnet when $K_{0}=0$ i.e., $Q=|{\bf Q}|=|D/2J|$ and/or $Q=0$ Nagaosa2012 . When $K_{0}\neq 0$, $\bf M$} and $\psi$ are correlated via the Kondo coupling. The energy spectrum of the conduction electron $\psi$ is shown in Fig. 1. The zero mode solutions are localized at the boundaries of the magnetic Brillouin zone which is introduced due to the helimagnet (for details, see Appendix A).
Figure 1: (color online) The energy spectrum in the continuum limit for Weyl fermions coupled with a helimagnet along the $q_{z}$-direction. We set $Q=K_{0}=1$. The helimagnet introduces a magnetic Brillouin zone. The red points stand for the Weyl nodes and the blue points stand for the zero modes (see Appendix A).
@@
and $\Phi=(c_{+\uparrow},c_{+\downarrow},c_{-\uparrow},c_{-\downarrow})$ with $\pm$ labeling two Weyl valleys; $\sigma$ and $\tau$ are the matrices that act on spin and valley indices, respectively. $k_{x}^{\prime}=2t_{x}^{\prime}k_{x}$, $k_{y}^{\prime}=2t_{y}^{\prime}k_{y}$, $k_{z}^{\prime}=2t_{z}(\sin{k_{0}})k_{z}$, $a_{0}={v_{0}(\cot{k_{0}})}/{t}$, and $b_{0}=2v_{0}(\sin{k_{0}})$. The spectra of Hamiltonian (4) read
$E_{\pm,\pm}=a_{0}k^{\prime}_{z}\pm$
$\sqrt{{\bf k}^{\prime 2}+(K_{0}/2)^{2}+b_{0}^{2}\pm\sqrt{K_{0}^{2}(k_{y}^{\prime 2}+k_{z}^{\prime 2})+4b_{0}^{2}{\bf k}^{\prime 2}}}.$ (5)
In the $K_{0}\rightarrow 0$ limit, $b_{0}$ becomes the energy offset between the two Weyl nodes, and $a_{0}$ controls the type of the Weyl node. When $b_{0}\rightarrow 0$, the CNHMC does not open a gap but drives the Weyl semimetal to a nodal-ring semimetal when $k_{y}^{\prime 2}+k_{z}^{\prime 2}=K_{0}^{2}/4$.
III Band structure and phase diagram
III.1 Band structures of the fermion
We now turn back to the lattice model. Similar to dealing with integer quantum Hall effects with the lattice model Kohmoto1989 ; Kohmoto1990 ; Hatsugai1993 (also see Appendix B for a brief review), one defines $|\Psi\rangle=\sum_{j=1}^{q}\psi_{j}c^{\dagger}(k_{x},k_{y},k_{z}^{0}+Qj)|0\rangle,$ where $Q=2\pi p/q$ with $p$ and $q$ coprime, $k_{z}=k_{z}^{0}+Qj$, and $\psi_{j+q}=\psi_{j}$. Then, the Schrödinger equation $H_{0}|\Psi\rangle=E|\Psi\rangle$ is reduced to the Harper equations
$[$ $m-2t_{x}\cos k_{x}-2t_{y}\cos k_{y}-2t_{z}\cos(k_{z}^{0}+Qj)]\sigma_{z}\psi_{j}$ (6)